# No [CII] or dust detection in two Little Red Dots at

### v1

@@ 3.2.1 SED fitting analysis
In this section, we compare the locations of our LRDs with the well-established –SFRtot scaling relation to investigate their alignment with expectations for star-forming galaxies or the potential influence of AGN activity. The –SFRtot relation has been shown to hold across a wide range of redshifts, with no significant evolution observed up to (e.g., Lagache et al., 2018; Schaerer et al., 2020; Carniani et al., 2020; Pallottini et al., 2022). Additionally, studies indicate that the relation is consistent for both AGN-hosting galaxies and typical star-forming galaxies, suggesting that the presence of AGN does not significantly alter the global –SFRtot relation (e.g., Herrera-Camus et al., 2018).
@@
For ID9094, when no AGN component is included in the SED fitting, the SFR derived corresponds to K. Combining this SFR with the 3 upper limits of the (log; Sect. 3.1) places ID9094 significantly below the –SFRtot relation, with a deviation of approximately dex (see Fig. 5). This large offset suggests that the observed properties of ID9094 cannot be explained solely by typical star formation processes in dusty star-forming galaxies. In contrast, when including an AGN component in the SED fitting, the resulting SFR with the K template is about four times lower than that returned by SED fitting without an AGN component. In this case, ID9094 lies within the scatter of the –SFRtot relation, indicating consistency with the scaling relation. This result strongly favors the AGN scenario for ID9094, providing the second hint of an AGN scenario.
@@
For ID2756, its 3 upper limits of the [C ii] emission line luminosity is log (Sect. 3.1). Due to the shallower NOEMA observations, we cannot constrain SFR well with and without AGN components in the SED (as described in Sect. 3.2.1). Therefore, we are unable to place meaningful constraints on its position in the versus SFRtot plane without deeper FIR observations.
In summary, the –SFRtot relation analysis provides a second hint in this work of AGN activity in the LRDs, complementing the first hint derived from the SED fitting analysis. These two independent methods, together with the broad H emission, consistently support the presence of an AGN in the two LRDs of our study, highlighting the critical role AGN-driven processes play in shaping their observed properties.
@@ 4 Impact of LRDs on our understanding of DSFGs
As introduced earlier, the two LRDs analyzed in this study were initially selected as the most massive DSFG candidates at in the JWST FRESCO GOODS-North field (Xiao et al., 2024). This misclassification arose due to the overlap in selection criteria between LRDs and DSFGs, particularly for optical/NIR-selected DSFGs – often referred to as optically dark/faint galaxies (also known as HST-dark galaxies or H-dropouts; e.g., Wang et al., 2019; Alcalde Pampliega et al., 2019; Williams et al., 2019; Xiao et al., 2023; Gómez-Guijarro et al., 2022; Barrufet et al., 2023; Akins et al., 2023; Pérez-González et al., 2023; Barro et al., 2024). Typically, DSFGs identified in optical/NIR observations are selected based on their red colors and rest-frame optical faintness, as these characteristics correlate strongly with dust attenuation at similar redshifts (e.g., Wang et al., 2016; Xiao et al., 2023). Additionally, the optical faintness criterion helps exclude quiescent galaxies, which are generally optically bright. For instance, DSFGs are often selected using a color cut such as F150W–F444W mag and F150W mag, though specific band choices and thresholds slightly vary across studies.
@@
With the increasing availability of JWST spectroscopic data, emission-line detections have become an additional criterion for DSFG selection (e.g., Xiao et al., 2024; Barrufet et al., 2025). In Xiao et al. (2024), DSFG candidates were selected based on their red color (F182M F444W ¿ 1.5 mag) and strong emission lines (e.g., H+[NII]+[SII] lines or [O iii]+H lines). This approach also allows for the identification of LRDs among DSFG candidates. Among 26 DSFG candidates with detected H lines (), seven () were found to be LRDs, exhibiting broad H emission lines (see Matthee et al., 2024).
@@
Over the past decade, numerous studies have shown that optically-dark/faint galaxies tend to have high stellar masses and may dominate the massive end of the stellar mass function (e.g., Wang et al., 2019; Gottumukkala et al., 2024). Additionally, these galaxies could contribute significantly to the cosmic star formation rate density (SFRD) at , with contributions of of the SFRD, depending on the methodology (see a direct comparison in Xiao et al. 2023; e.g., Wang et al. 2019; Williams et al. 2019; Gruppioni et al. 2020; Fudamoto et al. 2021; Talia et al. 2021; Enia et al. 2022; Shu et al. 2022; Barrufet et al. 2023; van der Vlugt et al. 2023). Notably, while different studies employ varying selection criteria for optically dark/faint galaxies, some adopt stricter definitions, requiring FIR/submm (e.g., ALMA/NOEMA) detection and/or non-detection in HST. Given that LRDs have been found to 1) lack FIR/submm detections (see also e.g., Labbe et al., 2023; Williams et al., 2024), 2) exhibit a characteristic “v-shaped” SED (indicating detection in HST), they are unlikely to contaminate the DSFG sample in these studies.
@@
On a global scale, these findings suggest that LRD contamination should not significantly bias SFRD estimates derived from DSFG samples. However, their presence may still impact number density estimates of optical/NIR-selected DSFG populations, particularly when broader color-based selection criteria are used. A more statistical study is needed to determine how much the number density estimates of DSFG populations are affected by LRDs. These results highlight the importance of refining DSFG selection methods to minimize potential contamination from AGN-dominated LRDs, ensuring a more accurate census of dust-obscured star formation in the early Universe.
@@ 5 Conclusions
In this paper, we present a study of two LRDs, ID9094 and ID2756, recently discovered in the JWST FRESCO GOODS-North field at . These sources, characterized by their compact, red morphology and ”v-shaped” SEDs, exhibit broad H emission lines, suggesting them as AGN candidates. Using NOEMA observations targeting the [C ii] emission line, and the 1.3 mm dust continuum, we obtain the following key results:
Both sources are undetected in the [C ii] and 1.3 mm continuum. For ID9094, the [C ii] and 1.3 mm upper limits are far below expectations for a typical SFG, where it would have been detected at in [C ii] and at 1.3 mm. For ID2756, the expected [C ii] detection at is also not observed.
Benefiting from the 1.3 mm upper limits, we conduct UV-to-FIR SED fitting, which provides the first hint of an AGN scenario in this work. For ID9094, we find that without an AGN component, the SED fitting suggests an unreasonably high dust temperature ( K), significantly exceeding typical values for SFGs at similar redshifts. Including an AGN component, however, yields more reasonable dust temperatures ( K), supporting the presence of AGN activity. The analysis of ID2756 is less conclusive due to the shallower NOEMA observations, underscoring the need for deeper FIR/submillimeter data.
Leveraging the [C ii] upper limits, we compare the sources to the –SFRtot relation, providing the second hint of an AGN scenario in this work. For ID9094, its [C ii] upper limit places it significantly below the scaling relation when no AGN component is included, but consistent with the relation when an AGN contribution is considered. This provides further evidence for the AGN scenario in ID9094, with implications for ID2756 pending further observations.
These results underscore the power of FIR observations, combined with short-wavelength photometry, in constraining the nature of LRDs, especially in the absence of spectroscopic data. For ID9094, the FIR constraints strongly support an AGN-dominated scenario, while the current data for ID2756 remain inconclusive, highlighting the need for deeper FIR/submillimeter observations. More broadly, this study demonstrates the critical role of FIR data in resolving the ambiguous classification of LRDs, which are often difficult to distinguish using optical or near-infrared photometry alone. Future deeper FIR/submillimeter campaigns, along with high-resolution spectroscopy and multi-wavelength data, will be essential to fully understand the physical properties of these LRDs and their role in early galaxy evolution.
@@ Acknowledgements.
This work is based on observations carried out under project number S23CY with the IRAM NOEMA Interferometer. IRAM is supported by INSU/CNRS (France), MPG (Germany) and IGN (Spain). This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs # 1895 and #4762. Support for this work was provided by NASA through grants JWST-GO-01895 and JWST-GO-04762 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-26555. This work has received funding from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00072, as well as from the Swiss National Science Foundation (SNSF) through project grant 200020_207349. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. This work is partially supported from the National Natural Science Foundation of China (12073003, 11721303, 11991052), and the China Manned Space Project (CMS-CSST-2021-A04 and CMS-CSST-2021-A06). Y.F. is supported by JSPS KAKENHI Grant Numbers JP22K21349 and JP23K13149. M.V. gratefully acknowledges financial support from the Independent Research Fund Denmark via grant numbers DFF 8021-00130 and 3103-00146 and from the Carlsberg Foundation via grant CF23-0417. VK acknowledges support from the University of Texas at Austin Cosmic Frontier Center. S.F. acknowledges support from NASA through the NASA Hubble Fellowship grant HST-HF2-51505.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. Support for this work for RPN was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51515.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555.References
- Akins et al. (2023) Akins, H. B., Casey, C. M., Allen, N., et al. 2023, ApJ, 956, 61
- Akins et al. (2025) Akins, H. B., Casey, C. M., Berg, D. A., et al. 2025, ApJ, 980, L29
- Akins et al. (2024) Akins, H. B., Casey, C. M., Lambrides, E., et al. 2024, arXiv e-prints, arXiv:2406.10341
- Alcalde Pampliega et al. (2019) Alcalde Pampliega, B., Pérez-González, P. G., Barro, G., et al. 2019, ApJ, 876, 135
- Ananna et al. (2024) Ananna, T. T., Bogdán, Á., Kovács, O. E., Natarajan, P., & Hickox, R. C. 2024, ApJ, 969, L18
- Baggen et al. (2024) Baggen, J. F. W., van Dokkum, P., Brammer, G., et al. 2024, ApJ, 977, L13
- Barro et al. (2024) Barro, G., Pérez-González, P. G., Kocevski, D. D., et al. 2024, ApJ, 963, 128
- Barrufet et al. (2025) Barrufet, L., Oesch, P. A., Marques-Chaves, R., et al. 2025, MNRAS[arXiv:2404.08052]
- Barrufet et al. (2023) Barrufet, L., Oesch, P. A., Weibel, A., et al. 2023, MNRAS, 522, 449
@@
- Boquien et al. (2019) Boquien, M., Burgarella, D., Roehlly, Y., et al. 2019, A&A, 622, A103
@@
- Calzetti et al. (2000) Calzetti, D., Armus, L., Bohlin, R. C., et al. 2000, ApJ, 533, 682
- Carniani et al. (2020) Carniani, S., Ferrara, A., Maiolino, R., et al. 2020, MNRAS, 499, 5136
- Casey (2012) Casey, C. M. 2012, MNRAS, 425, 3094
- Casey et al. (2024) Casey, C. M., Akins, H. B., Shuntov, M., et al. 2024, ApJ, 965, 98
@@
- Cowie et al. (2017) Cowie, L. L., Barger, A. J., Hsu, L. Y., et al. 2017, ApJ, 837, 139
- da Cunha et al. (2013) da Cunha, E., Groves, B., Walter, F., et al. 2013, ApJ, 766, 13
- De Looze et al. (2014) De Looze, I., Cormier, D., Lebouteiller, V., et al. 2014, A&A, 568, A62
- De Rossi et al. (2018) De Rossi, M. E., Rieke, G. H., Shivaei, I., Bromm, V., & Lyu, J. 2018, ApJ, 869, 4
- Decarli et al. (2023) Decarli, R., Pensabene, A., Diaz-Santos, T., et al. 2023, A&A, 673, A157
- D’Eugenio et al. (2024) D’Eugenio, F., Cameron, A. J., Scholtz, J., et al. 2024, arXiv e-prints, arXiv:2404.06531
- Egami et al. (2023) Egami, E., Sun, F., Alberts, S., et al. 2023, Complete NIRCam Grism Redshift Survey (CONGRESS), JWST Proposal. Cycle 2, ID. #3577
- Eisenstein et al. (2023) Eisenstein, D. J., Willott, C., Alberts, S., et al. 2023, arXiv e-prints, arXiv:2306.02465
- Elbaz et al. (2011) Elbaz, D., Dickinson, M., Hwang, H. S., et al. 2011, A&A, 533, A119
- Endsley et al. (2022) Endsley, R., Stark, D. P., Bouwens, R. J., et al. 2022, MNRAS, 517, 5642
- Enia et al. (2022) Enia, A., Talia, M., Pozzi, F., et al. 2022, ApJ, 927, 204
@@
- Fritz et al. (2006) Fritz, J., Franceschini, A., & Hatziminaoglou, E. 2006, MNRAS, 366, 767
- Fudamoto et al. (2024) Fudamoto, Y., Inoue, A. K., Coe, D., et al. 2024, ApJ, 961, 71
- Fudamoto et al. (2021) Fudamoto, Y., Oesch, P. A., Schouws, S., et al. 2021, Nature, 597, 489
- Fujimoto et al. (2022) Fujimoto, S., Brammer, G. B., Watson, D., et al. 2022, Nature, 604, 261
- Gardner et al. (2023) Gardner, J. P., Mather, J. C., Abbott, R., et al. 2023, PASP, 135, 068001
@@
- Gottumukkala et al. (2024) Gottumukkala, R., Barrufet, L., Oesch, P. A., et al. 2024, MNRAS, 530, 966
- Greene et al. (2024) Greene, J. E., Labbe, I., Goulding, A. D., et al. 2024, ApJ, 964, 39

### v2

@@ 3.2.1 SED fitting analysis
In this section, we compare the locations of our LRDs with the well-established āSFRtot scaling relation to investigate their alignment with expectations for star-forming galaxies or the potential influence of AGN activity. The āSFRtot relation has been shown to hold across a wide range of redshifts, with no significant evolution observed up to (e.g., Lagache etĀ al., 2018; Schaerer etĀ al., 2020; Carniani etĀ al., 2020; Pallottini etĀ al., 2022). Additionally, studies indicate that the relation is consistent for both AGN-hosting galaxies and typical star-forming galaxies, suggesting that the presence of AGN does not significantly alter the global āSFRtot relation (e.g., Herrera-Camus etĀ al., 2018).
@@
For ID9094, when no AGN component is included in the SED fitting, the SFR derived corresponds to K. Combining this SFR with the 3 upper limits of the (log(; Sect.Ā 3.1) places ID9094 significantly below the āSFRtot relation, with a deviation of dex (see Fig.Ā 5). This large offset suggests that the observed properties of ID9094 may not be fully explained by typical star formation processes in dusty star-forming galaxies. In contrast, when including an AGN component in the SED fitting, the resulting SFR with the K template is about four times lower than that returned by SED fitting without an AGN component. In this case, ID9094 lies within the scatter of the āSFRtot relation, indicating consistency with the scaling relation. We note that including an AGN component with K results in a lower SFR and brings ID9094 into even better agreement with the āSFRtot relation.
@@
For ID2756, its 3 upper limits of the [Cāii]ā emission line luminosity is log( (Sect.Ā 3.1). Due to the shallower NOEMA observations, we cannot constrain SFR well with and without AGN components in the SED (as described in Sect.Ā 3.2.1). Therefore, we are unable to place meaningful constraints on its position in the versus SFRtot plane without deeper FIR observations.
3.2.3 Caveats
Our analysis provides constraints on the dust and gas properties of the two LRDs. For ID9094, the two independent analyses of and the āSFRtot relation show notable deviations from typical dusty star-forming galaxies. These distinct properties suggest that additional radiation mechanisms, such as AGN activity, may contribute to the total emission of this source.
However, we emphasize that the interpretations above rely on several assumptions and modeling choices. First, the lower limit on the dust temperature (āK) inferred in the absence of an AGN component is derived under fixed SED parameters and based solely on upper limits in the FIR. The resulting value is therefore model-dependent and sensitive to assumptions such as the dust emissivity index and IR SED shape. Moreover, high dust temperatures alone are not sufficient to definitively rule out a purely star-forming scenario. Although rare, some high-redshift galaxies without clear AGN signatures have exhibited elevated values. For instance, MACS0416-Y1 at has a K, potentially due to intense starburst activity in a merger-driven system without requiring AGN heating (e.g., Bakx etĀ al., 2020; Fudamoto etĀ al., 2023; Harshan etĀ al., 2024; Sommovigo etĀ al., 2022). Similarly, A2744-YD4 at shows K, which may be attributed to its low metallicity and high star formation rate (e.g., Laporte etĀ al., 2017; Behrens etĀ al., 2018; Sommovigo etĀ al., 2022). These cases are considered extreme and likely involve compact starbursts and low dust content.
Second, the observed offset from the empirical āSFR relation is not conclusive evidence against the DSFG scenario. The total SFR estimates are subject to uncertainties, especially when derived from upper limits, and the comparison sample includes a mix of UV-bright and IR-selected galaxies with diverse properties. Moreover, the āSFR relation exhibits large scatter at high redshift, and many star-forming galaxies also lie below the relation.
In summary, while our results are more consistent with an AGN-driven scenario, the possibility of dusty star formation remains viable given the current data limitations. Future deeper FIR observations will be essential to distinguish between these scenarios.
@@ 4 Impact of LRDs on our understanding of DSFGs
As introduced earlier, the two LRDs analyzed in this study were initially selected as the most massive DSFG candidates at in the JWST FRESCO GOODS-North field (Xiao etĀ al., 2024). This misclassification arose due to the overlap in selection criteria between LRDs and DSFGs, particularly for optical/NIR-selected DSFGs ā often referred to as optically dark/faint galaxies (also known as HST-dark galaxies or H-dropouts; e.g., Wang etĀ al., 2019; Alcalde Pampliega etĀ al., 2019; Williams etĀ al., 2019; Xiao etĀ al., 2023; Gómez-Guijarro etĀ al., 2022; Barrufet etĀ al., 2023; Akins etĀ al., 2023; PĆ©rez-GonzĆ”lez etĀ al., 2023; Barro etĀ al., 2024). Typically, DSFGs identified in optical/NIR observations are selected based on their red colors and rest-frame optical faintness, as these characteristics correlate strongly with dust attenuation at similar redshifts (e.g., Wang etĀ al., 2016; Xiao etĀ al., 2023). Additionally, the optical faintness criterion helps exclude quiescent galaxies, which are generally optically bright. For instance, DSFGs are often selected using a color cut such as F150WāF444W mag and F150W mag, though specific band choices and thresholds slightly vary across studies.
@@
With the increasing availability of JWST spectroscopic data, emission-line detections have become an additional criterion for DSFG selection (e.g., Xiao etĀ al., 2024; Barrufet etĀ al., 2025). In Xiao etĀ al. (2024), DSFG candidates were selected based on their red color (F182M F444W Āæ 1.5 mag) and strong emission lines (e.g., H+[NII]+[SII] lines or [Oāiii]+H lines). This approach also allows for the identification of LRDs among DSFG candidates. Among 26 DSFG candidates with detected H lines (), seven () were found to be LRDs, exhibiting broad H emission lines (see Matthee etĀ al., 2024).
@@
Over the past decade, numerous studies have shown that optically-dark/faint galaxies tend to have high stellar masses and may dominate the massive end of the stellar mass function (e.g., Wang etĀ al., 2019; Gottumukkala etĀ al., 2024). Additionally, these galaxies could contribute significantly to the cosmic star formation rate density (SFRD) at , with contributions of of the SFRD, depending on the methodology (see a direct comparison in Xiao etĀ al. 2023; e.g., Wang etĀ al. 2019; Williams etĀ al. 2019; Gruppioni etĀ al. 2020; Fudamoto etĀ al. 2021; Talia etĀ al. 2021; Enia etĀ al. 2022; Shu etĀ al. 2022; Barrufet etĀ al. 2023; van der Vlugt etĀ al. 2023). Notably, while different studies employ varying selection criteria for optically dark/faint galaxies, some adopt stricter definitions, requiring FIR/submm (e.g., ALMA/NOEMA) detection and/or non-detection in HST. Given that LRDs have been found to 1) lack FIR/submm detections (see also e.g., Labbe etĀ al., 2025; Williams etĀ al., 2024), 2) exhibit a characteristic āv-shapedā SED (indicating detection in HST), they are unlikely to contaminate the DSFG sample in these studies.
@@
On a global scale, these findings suggest that LRD contamination should not significantly bias SFRD estimates derived from DSFG samples. However, their presence may still impact number density estimates of optical/NIR-selected DSFG populations, particularly when broader color-based selection criteria are used. A study with a larger sample is needed to determine how much the number density estimates of DSFG populations are affected by LRDs. These results highlight the importance of refining DSFG selection methods to minimize potential contamination from AGN-dominated LRDs, ensuring a more accurate census of dust-obscured star formation in the early Universe.
@@ 5 Conclusions
In this paper, we present a study of two LRDs, ID9094 and ID2756, recently discovered in the JWST FRESCO GOODS-North field at . These sources, characterized by their red, point-source-like morphology and āv-shapedā SEDs, exhibit broad H emission lines, suggesting them as AGN candidates. Using NOEMA observations targeting the [Cāii] emission line, and the 1.3 mm dust continuum, we obtain the following key results:
ā¢
Both sources are undetected in the [Cāii]Ā and 1.3 mm continuum. For ID9094, the upper limits on [Cāii]Ā and 1.3 mm continuum are far below expectations for a typical DSFG, where it would have been detected at in [Cāii]Ā and at 1.3 mm. For ID2756, the expected detection in [Cāii]Ā is also not observed.
ā¢
Benefiting from the 1.3 mm upper limits, we perform UV-to-FIR SED fitting, which provides the first hint of an AGN scenario in this work. For ID9094, without including an AGN component, the SED fit requires an abnormally high dust temperature ( K), significantly above typical values for SFGs at similar redshifts. Including an AGN component yields more physically plausible fits with dust temperatures that can be as low as 40ā60 K. However, we caution that the minimum is strongly linked to the AGN fraction contributing to the total IR luminosity ā higher AGN fractions allow for lower values. For ID2756, the SED fitting remains inconclusive due to shallower NOEMA data, highlighting the need for deeper FIR/submillimeter observations.
ā¢
Using the [Cāii]Ā upper limits, we compare the sources to the empirical āSFRtot relation, providing a second hint of an AGN scenario in this work. For ID9094, its [Cāii]Ā upper limit places it significantly below the scaling relation when no AGN component is included, but consistent with the relation when an AGN contribution is considered. The case for ID2756 remains uncertain due to limited NOEMA depth.
These results highlight the value of FIR constraints, when combined with shorter-wavelength photometry, in revealing the nature of LRDs ā particularly when spectroscopic data are unavailable. For ID9094, the results are consistent with a scenario in which AGN activity contributes to the observed emission, though a dusty star-forming origin cannot be fully excluded given the data limitations. We also note that both sources are detected in JWST/MIRI F1000W, and ID9094 is additionally detected in F2100W, supporting the presence of hot dust emission (G. Leung, priv. comm.). For ID2756, the current observations are insufficient to determine the dominant energy source. We emphasize that our interpretation remains model-dependent and subject to assumptions in the SED fitting and limited FIR constraints. More broadly, this study demonstrates the critical role of FIR data in clarifying the nature of LRDs. Future deeper FIR/submillimeter campaigns, along with high-resolution spectroscopy and multi-wavelength data, will be essential to fully understand the physical properties of these LRDs and their role in early galaxy evolution.
@@ Acknowledgements.
We are very grateful to the anonymous referee for instructive comments, which helped improve the overall quality and strengthen the analysis of this work. We thank Andrea Weibel for assistance with the HST and JWST photometric measurements used in this paper. This work is based on observations carried out under project number S23CY with the IRAM NOEMA Interferometer. IRAM is supported by INSU/CNRS (France), MPG (Germany) and IGN (Spain). This work is based in part on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1895 and #4762. Support for programs #1895 and #4762 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. This work has received funding from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00072, as well as from the Swiss National Science Foundation (SNSF) through project grant 200020_207349. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. This work is partially supported from the National Natural Science Foundation of China (12073003, 11721303, 11991052), and the China Manned Space Project (CMS-CSST-2021-A04 and CMS-CSST-2021-A06). Y.F.Ā is supported by JSPS KAKENHI Grant Numbers JP22K21349 and JP23K13149. M.V.Ā gratefully acknowledges financial support from the Independent Research Fund Denmark via grant numbers DFF 8021-00130 and 3103-00146 and from the Carlsberg Foundation via grant CF23-0417. VK acknowledges support from the University of Texas at Austin Cosmic Frontier Center. S.F. acknowledges support from NASA through the NASA Hubble Fellowship grant HST-HF2-51505.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. Support for this work for RPN was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51515.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555.References
- Akins etĀ al. (2023) Akins, H.Ā B., Casey, C.Ā M., Allen, N., etĀ al. 2023, ApJ, 956, 61
- Akins etĀ al. (2025a) Akins, H.Ā B., Casey, C.Ā M., Berg, D.Ā A., etĀ al. 2025a, ApJ, 980, L29
- Akins etĀ al. (2025b) Akins, H.Ā B., Casey, C.Ā M., Chisholm, J., etĀ al. 2025b, submitted to ApJL, arXiv:2503.00998
- Akins etĀ al. (2024) Akins, H.Ā B., Casey, C.Ā M., Lambrides, E., etĀ al. 2024, submitted to ApJ, arXiv:2406.10341
- Alcalde Pampliega et al. (2019) Alcalde Pampliega, B., Pérez-GonzÔlez, P. G., Barro, G., et al. 2019, ApJ, 876, 135
- Ananna etĀ al. (2024) Ananna, T.Ā T., BogdĆ”n, Ć., KovĆ”cs, O.Ā E., Natarajan, P., & Hickox, R.Ā C. 2024, ApJ, 969, L18
- Baggen etĀ al. (2024) Baggen, J. F.Ā W., van Dokkum, P., Brammer, G., etĀ al. 2024, ApJ, 977, L13
- Bakx etĀ al. (2020) Bakx, T. J.Ā L.Ā C., Tamura, Y., Hashimoto, T., etĀ al. 2020, MNRAS, 493, 4294
- Barro et al. (2024) Barro, G., Pérez-GonzÔlez, P. G., Kocevski, D. D., et al. 2024, ApJ, 963, 128
- Barrufet etĀ al. (2025) Barrufet, L., Oesch, P.Ā A., Marques-Chaves, R., etĀ al. 2025, MNRAS, 537, 3453
- Barrufet etĀ al. (2023) Barrufet, L., Oesch, P.Ā A., Weibel, A., etĀ al. 2023, MNRAS, 522, 449
- Behrens etĀ al. (2018) Behrens, C., Pallottini, A., Ferrara, A., Gallerani, S., & Vallini, L. 2018, MNRAS, 477, 552
@@
- Boquien etĀ al. (2019) Boquien, M., Burgarella, D., Roehlly, Y., etĀ al. 2019, A&A, 622, A103
@@
- Calzetti etĀ al. (2000) Calzetti, D., Armus, L., Bohlin, R.Ā C., etĀ al. 2000, ApJ, 533, 682
- Carniani etĀ al. (2020) Carniani, S., Ferrara, A., Maiolino, R., etĀ al. 2020, MNRAS, 499, 5136
- Casey (2012) Casey, C.Ā M. 2012, MNRAS, 425, 3094
- Casey etĀ al. (2025) Casey, C.Ā M., Akins, H.Ā B., Finkelstein, S.Ā L., etĀ al. 2025, submitted to AAS journals, arXiv:2505.18873
- Casey etĀ al. (2024) Casey, C.Ā M., Akins, H.Ā B., Shuntov, M., etĀ al. 2024, ApJ, 965, 98
@@
- Cowie etĀ al. (2017) Cowie, L.Ā L., Barger, A.Ā J., Hsu, L.Ā Y., etĀ al. 2017, ApJ, 837, 139
- da Cunha etĀ al. (2013) da Cunha, E., Groves, B., Walter, F., etĀ al. 2013, ApJ, 766, 13
- de Graaff etĀ al. (2025) de Graaff, A., Rix, H.-W., Naidu, R.Ā P., etĀ al. 2025, submitted to A&A, arXiv:2503.16600
- De Looze etĀ al. (2014) De Looze, I., Cormier, D., Lebouteiller, V., etĀ al. 2014, A&A, 568, A62
- De Rossi etĀ al. (2018) De Rossi, M.Ā E., Rieke, G.Ā H., Shivaei, I., Bromm, V., & Lyu, J. 2018, ApJ, 869, 4
- Decarli etĀ al. (2023) Decarli, R., Pensabene, A., Diaz-Santos, T., etĀ al. 2023, A&A, 673, A157
- DāEugenio etĀ al. (2025) DāEugenio, F., Cameron, A.Ā J., Scholtz, J., etĀ al. 2025, ApJS, 277, 4
- Egami etĀ al. (2023) Egami, E., Sun, F., Alberts, S., etĀ al. 2023, Complete NIRCam Grism Redshift Survey (CONGRESS), JWST Proposal. Cycle 2, ID. #3577
- Eisenstein etĀ al. (2023) Eisenstein, D.Ā J., Willott, C., Alberts, S., etĀ al. 2023, submitted to ApJS, arXiv:2306.02465
- Elbaz etĀ al. (2011) Elbaz, D., Dickinson, M., Hwang, H.Ā S., etĀ al. 2011, A&A, 533, A119
- Endsley etĀ al. (2022) Endsley, R., Stark, D.Ā P., Bouwens, R.Ā J., etĀ al. 2022, MNRAS, 517, 5642
- Enia etĀ al. (2022) Enia, A., Talia, M., Pozzi, F., etĀ al. 2022, ApJ, 927, 204
@@
- Fritz etĀ al. (2006) Fritz, J., Franceschini, A., & Hatziminaoglou, E. 2006, MNRAS, 366, 767
- Fudamoto etĀ al. (2024) Fudamoto, Y., Inoue, A.Ā K., Coe, D., etĀ al. 2024, ApJ, 961, 71
- Fudamoto etĀ al. (2023) Fudamoto, Y., Inoue, A.Ā K., & Sugahara, Y. 2023, MNRAS, 521, 2962
- Fudamoto etĀ al. (2021) Fudamoto, Y., Oesch, P.Ā A., Schouws, S., etĀ al. 2021, Nature, 597, 489
- Fujimoto etĀ al. (2022) Fujimoto, S., Brammer, G.Ā B., Watson, D., etĀ al. 2022, Nature, 604, 261
- Gardner etĀ al. (2023) Gardner, J.Ā P., Mather, J.Ā C., Abbott, R., etĀ al. 2023, PASP, 135, 068001
@@
- Gordon etĀ al. (2010) Gordon, K.Ā D., Galliano, F., Hony, S., etĀ al. 2010, A&A, 518, L89
- Gottumukkala etĀ al. (2024) Gottumukkala, R., Barrufet, L., Oesch, P.Ā A., etĀ al. 2024, MNRAS, 530, 966
- Greene etĀ al. (2024) Greene, J.Ā E., Labbe, I., Goulding, A.Ā D., etĀ al. 2024, ApJ, 964, 39